# Ultralight Black Holes as Sources of High-Energy Particles

### v3

@@ Abstract
Results — Figure 1 shows the integrated gamma-ray flux $\Phi_{i}^{\rm gal}(E)$ in Eq. (6), produced by newly-formed BHs from PBH mergers and for $f_{\rm PBH}=1$, for different values of the PBH masses and the conservative merger rate in Eq. (5). We have not included the eg contribution since, at energies above the TeV scale, the Universe is opaque to gamma rays due to gamma-gamma absorption (see Cao et al. (2023) for a review). Results are compared with the upper limits at 95% confidence level (CL) from the Pierre Auger Observatory (Auger) Savina (2021); Abreu et al. (2022, 2023), the Karlsruhe Shower Core and Array Detector (KASCADE), and KASCADE-Grande Apel et al. (2017). Lighter BHs are brighter due to the higher Hawking temperature and smaller entropy. For this, only the heavier BHs evade the bounds from the measured flux of gamma rays. We have not included any effects from physics beyond the SM, which would impact our results by altering the semiclassical evaporation timescale and the emission from the secondary spectrum.
@@
Figure 1: The integrated gamma-ray flux emitted by newly-formed BHs from memory-burdened Galactic PBH mergers, for a monochromatic PBHs distribution with $f_{\rm PBH}=1$. Also shown are the upper limits at 95% CL from Auger Savina (2021); Abreu et al. (2022, 2023), KASCADE, and KASCADE-Grande Apel et al. (2017).
@@
We now discuss the emission of energetic neutrinos of all flavors by the same sources. The differential neutrino flux per solid angle and energy is reported by the IceCube collaboration using the cascades collected over 7.5 years of operation at energies above 60 TeV Abbasi et al. (2021a); Aartsen et al. (2020). Above the PeV, an upper limit at 95% CL on the neutrino flux is reported Aartsen et al. (2018, 2019). In Fig. 2 we compare these experimental results with the the differential flux of all-flavor neutrinos sourced by newly-formed BHs. We include the Galactic (continuous) and extragalactic (dashed) contributions assuming the merger rate in Eq. (5).
Figure 2: The differential flux of all-flavor neutrinos sourced by newly-formed BHs, for a monochromatic PBHs distribution with $f_{\rm PBH}=1$. We show the Galactic (continuous) and the extra-galactic (dashed) contributions. The theoretical flux is compared with the all-flavor neutrino flux above 60 TeV Abbasi et al. (2021a); Aartsen et al. (2020) and with the upper bound for ultra-high energy neutrinos Aartsen et al. (2018, 2019).
@@
Discussion — Assuming that a fraction $f_{\rm PBH}$ of the DM relic abundance consists of light PBHs, we derive a constraint on the model in the $(m_{\rm PBH},f_{\rm PBH})$ plane by performing a least-square analysis that accounts for the neutrino flux measurements and upper bounds reported by the IceCube collaboration Abbasi et al. (2021a); Aartsen et al. (2020, 2018, 2019). Using the merger rate in Eq. (5), Fig. 3 excludes light PBHs of masses $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$ as the totality of the DM. The mass range is further constrained, with the upper bound $m_{\rm PBH}\lesssim 10^{10}$ g from considerations over the effects of PBH evaporation at the onset of Big-Bang nucleosynthesis (BBN, blue contour) Carr et al. (2010). Also shown is the value $m_{\rm PBH}\sim 3\times 10^{3}$ g, corresponding to PBHs evaporating today when $k=1$ in Eq. (2) (vertical dashed line). For $k\lesssim 1$, constraints have been placed on PBH masses up to approximately $m_{\rm PBH}\sim 10^{5}$ g from the emission of photons Alexandre et al. (2024); Thoss et al. (2024) or neutrinos Chianese et al. (2024). However, the signal following the merger considered here is independent of $k$, offering a novel and complementary signature that does not depend on the specific model building details of the stabilized phase. Our results suggest that the window where memory-burdened PBHs constitute the DM is in tension with current measurements on the neutrino flux. However, this enhancement stems from limited theoretical understanding of non-Gaussian effects in PBH clustering and tidal disruption of PBH halos Delos et al. (2024). The bounds would be relaxed if the memory burden effect sets in earlier than the timescale $\tau_{\rm sc}$, resulting in a larger remnant mass Dvali et al. (2020, 2024a).
@@
Figure 3: The DM fraction in PBHs of mass $m_{\rm PBH}$ as a function of the PBH mass. The results in this work consist in the bound from IceCube. Also shown are bounds from BBN considerations. In deriving the bound, we assumed that the memory burden effect occurs when the PBH has lost half of its initial mass. Different mass loss values are possible, potentially altering our findings as discussed in the text.
@@
The results in Fig. 3 suggest that PBHs could make up a sizeable fraction of the DM in two mass ranges: $m_{\rm PBH}\lesssim 10^{5}$ g and $10^{8}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{9}$. The lower mass range can be probed through gamma-ray flux measurements from experiments such as Auger (light green) Aab et al. (2020); Abreu et al. (2021a, b), Telescope Array (cyan) Abu-Zayyad et al. (2013); Ivanov (2020), and the Large High Altitude Air Shower Observatory (LHAASO) Addazi et al. (2022), which can detect gamma-rays above about 50 TeV. Studies of gamma and cosmic ray emissions from PBH clusters could further constraints $f_{\rm PBH}$, particularly for masses between $10^{3}\textrm{--}10^{5}$ g, while cosmic ray anisotropies might reveal nearby PBH subhalos Abbasi et al. (2010, 2011); Abeysekara et al. (2014, 2019); Amenomori et al. (2021), though more data is needed Giacinti and Sigl (2012). For higher-mass PBHs, upcoming neutrino observatories like IceCube-Gen2 radio array Abbasi et al. (2021b), the Radio Neutrino Observatory in Greenland (RNO-G) Aguilar et al. (2021), and the Giant Radio Array for Neutrino Detection (GRAND) Álvarez-Muñiz et al. (2020), will probe neutrinos at energies above 10 PeV with improved sensitivity. These observatories could constrain $f_{\rm PBH}$ by analyzing neutrino fluxes, especially when combined with data from Galactic sources, improving detection reach by $\mathcal{O}(10)$ over 10 years. Directional neutrino measurements will be crucial for distinguishing PBH contributions, providing a test for this scenario. Evaporating PBHs may emit gravitons, peaking at frequencies $\gtrsim 10^{30}$ Hz for $m_{\rm PBH}\lesssim 10^{8}$ g, beyond the reach of current GW detectors Domcke et al. (2022); Schmieden and Schott (2024); Gatti et al. (2024). GW abundance estimates suggest $\Omega_{\rm GW}\lesssim 10^{-12}$ at the peak.
@@
Conclusions — In this Letter, we have explored the astrophysical implications to a PBH population affected by the memory burden. Given the halt in the evaporation process, light PBHs might live over cosmological timescales as the DM, undergoing binary merging and forming new BHs of similar mass at much smaller redshift than their formation epochs. We argued that newly-formed BHs initially do not carry memory burden and would then emit energetic particles in the form of cosmic rays, gamma rays, and neutrinos, which can be searched at dedicated facilities. We have estimated the integral spectrum in gamma rays from Galactic mergers and the differential spectrum of all-flavor neutrinos. Based on the current measurements by IceCube, we have derived the bounds on the PBH fraction, showing the tension with obtaining PBHs as the DM in the mass range $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$. If the merger rate is increased by primordial non-Gaussianity, the mass range that is potentially excluded sensibly widens. Future detectors will easily reach this signal. Other future work in this direction includes the careful study of the model with cosmic-rays from Galactic substructures.
@@ Acknowledgements.
We thank Gia Dvali for useful discussions and for reading the manuscript. We further thank Gwenael Giacinti for useful comments. Finally, we are grateful to Concha Gonzalez-Garcia for correcting a previous mistake in the detectability of the neutrino mass nature at IceCube. We acknowledge support by the National Natural Science Foundation of China (NSFC) through the grant No. 12350610240 “Astrophysical Axion Laboratories”, as well as hospitality by the Istituto Nazionale di Fisica Nucleare (INFN).

### v4

@@ III Results
Figure 1 shows the integrated gamma-ray flux $\Phi_{i}^{\rm gal}(E)$ in Eq. (6), produced by newly-formed BHs from PBH mergers and for $f_{\rm PBH}=1$, for different values of the PBH masses and the conservative merger rate in Eq. (5). The absorption of gamma rays over cosmological timescales, as well as cascade effects, are implemented following Ref. Dondarini et al. (2025). Results are compared with the upper limits at 95% confidence level (CL) from the Pierre Auger Observatory (Auger) Savina (2021); Abreu et al. (2022, 2023), the Karlsruhe Shower Core and Array Detector (KASCADE), and KASCADE-Grande Apel et al. (2017). Lighter BHs are brighter due to the higher Hawking temperature and smaller entropy. For this, only the heavier BHs evade the bounds from the measured flux of gamma rays. We have not included any effects from physics beyond the SM, which would impact our results by altering the semiclassical evaporation timescale and the emission from the secondary spectrum.
@@
Figure 1: The integrated gamma-ray flux emitted by newly-formed BHs from memory-burdened Galactic PBH mergers, for a monochromatic PBHs distribution with $f_{\rm PBH}=1$. $m_{i}$ denotes the mass of the PBHs post-merger which is roughly twice the mass of the DM. Also shown are the upper limits at 95% CL from Auger Savina (2021); Abreu et al. (2022, 2023), KASCADE, and KASCADE-Grande Apel et al. (2017).
@@
We now discuss the emission of energetic neutrinos by the same sources. The differential neutrino flux per solid angle and energy is reported by the IceCube collaboration using the cascades collected over 7.5 years of operation at energies above 60 TeV Abbasi et al. (2021a); Aartsen et al. (2020). Below around 10 PeV, the neutrino flux is bound by observations from ANTARES Albert et al. (2024). At energies above 10 PeV, an upper limit at 95% CL on the neutrino flux is reported by the IceCube High-Energy Starting Events (HESE) Aartsen et al. (2018, 2019) and by Auger Abdul Halim et al. (2023). IceCube reports candidate detections through the Northern Sky Tracks (NST) Abbasi et al. (2022) and HESE Abbasi et al. (2021a). In Fig. 2 we show these measurements, together with the recent high-energy event detected by KM3NeT collaboration Aiello et al. (2025) as well as the Glashow resonant observed by IceCube Aartsen et al. (2021). We compare these experimental results with the per-flavor neutrinos differential flux sourced by merged PBHs. The Galactic (dashed), extragalactic (dot-dashed), and total (solid) contributions are obtained assuming the merger rate in Eq. (5).
Figure 2: Per-flavor neutrino fluxes for various monochromatic PBH mass distributions with $f_{\rm PBH}=1$, normalized by a factor of 2 22 for better readability. Dashed (dot-dashed) lines represent the Galactic (extra-galactic) contributions, while solid lines show their sum. The black line in the upper right corner indicates the current upper limit from IceCube High-Energy Starting Events (HESE) Aartsen et al. (2018), while the one on the left shows limits from ANTARES Albert et al. (2024). The gray line corresponds to the upper bound set by Auger Abdul Halim et al. (2023). Black (blue) points indicate IceCube detections from Northern Sky Tracks (NST) Abbasi et al. (2022) and HESE Abbasi et al. (2021a), respectively. The green cross marks the Glashow resonance event observed by IceCube Aartsen et al. (2021), and the red point shows the high-energy neutrino KM3-230213A detected by KM3NeT Aiello et al. (2025).
IV Discussion
Assuming that a fraction $f_{\rm PBH}$ of the DM relic abundance consists of light PBHs, we derive a constraint on the model in the $(m_{\rm PBH},f_{\rm PBH})$ plane by performing a least-square analysis that accounts for the neutrino flux measurements and upper bounds reported by the IceCube collaboration Abbasi et al. (2021a); Aartsen et al. (2020, 2018, 2019).
@@ IV Discussion
Using the merger rate in Eq. (5), Fig. 3 excludes light PBHs of masses $m_{\rm PBH}\lesssim 10^{9}\,\rm g$ as the totality of the DM (filled cyan area). The mass range is further constrained, with the upper bound $m_{\rm PBH}\lesssim 10^{10}$ g from considerations over the effects of PBH evaporation at the onset of Big-Bang nucleosynthesis (BBN, blue contour) Carr et al. (2010, 2021c); Thoss and Burkert (2025). Also shown is the value $m_{\rm PBH}\sim 3\times 10^{3}$ g (vertical dashed line), for which $\tau_{\rm sc}\simeq t_{0}$ when $k=2$, c.f., Eq. (2). For $k=2$, constraints have been placed on PBH masses up to approximately $m_{\rm PBH}\sim 10^{5}$ g from the emission of photons Alexandre et al. (2024); Thoss et al. (2024) or neutrinos Chianese et al. (2025) (see Dondarini et al. (2025) for a complete analysis). However, the signal following the merger considered here is not sensitive to $k$, offering a novel and complementary signature that does not depend on the specific details of the stabilized phase.
Our results suggest that the window where memory-burdened PBHs constitute the DM is in tension with current measurements on the neutrino flux.
@@
Figure 3: The DM fraction in PBHs of mass $m_{\rm PBH}$ as a function of the PBH mass. The results in this work consist in the bound from IceCube (filled cyan area). Also shown are bounds from BBN considerations, taken from Carr et al. (2010, 2021c); Thoss and Burkert (2025). In deriving the bound, we assume that the memory burden effect occurs when the PBH has lost half of its initial mass. Different mass-loss fraction are possible, potentially altering our findings as discussed in the text.
As shown in Fig. 3, only PBHs in the mass range $10^{9}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{10}$ could make up a sizable fraction of the DM. Obviously, it is possible that memory burden sets in before half of the mass has been emitted semi-classically. This possibility is actually supported in theoretical studies, where it has been found that memory burden may emerge after an $\mathcal{O}(1/\sqrt{S})$ fraction of the initial mass has been emitted Dvali (2018a); Dvali et al. (2019, 2020); Alexandre et al. (2024); Dvali et al. (2024a). Independently motivated studies found that, on comparable timescales, the semiclassical description of BHs ceases to be accurate Dvali and Gomez (2013a, b, 2014a, 2012); Dvali et al. (2013); Dvali and Panchenko (2015); Michel and Zell (2023). In such scenario, the bounds presented here would no longer apply.
@@
Comparable, and potentially stronger constraints are expected in a similar mass range due to gamma-ray flux measurements by the Large High Altitude Air Shower Observatory (LHAASO) Addazi et al. (2022), which detected gamma-rays above about 10 TeV. Studies of gamma and cosmic ray emissions from PBH clusters could further constraints $f_{\rm PBH}$, particularly for masses smaller than $10^{5}$ g, while cosmic ray anisotropies might reveal nearby PBH subhalos Abbasi et al. (2010, 2011); Abeysekara et al. (2014, 2019); Amenomori et al. (2021), though more data is needed Giacinti and Sigl (2012). For higher-mass PBHs, upcoming neutrino observatories like IceCube-Gen2 radio array Abbasi et al. (2021b), the Radio Neutrino Observatory in Greenland (RNO-G) Aguilar et al. (2021), and the Giant Radio Array for Neutrino Detection (GRAND) Álvarez-Muñiz et al. (2020), will probe neutrinos at energies above 10 PeV with improved sensitivity. In this direction, we ought to mention the recent measurement of the high-energy - above $10^{2}\,$eV - neutrino event detected by the KM3NeT collaboration Aiello et al. (2025). These observatories could constrain $f_{\rm PBH}$ by analyzing neutrino fluxes, especially when combined with data from Galactic sources, improving detection reach by $\mathcal{O}(10)$ over 10 years. Directional neutrino measurements will be crucial for distinguishing PBH contributions, providing a test for this scenario. Evaporating PBHs may emit gravitons, peaking at frequencies $\gtrsim 10^{30}$ Hz for $m_{\rm PBH}\lesssim 10^{8}$ g, beyond the reach of current GW detectors Domcke et al. (2022); Schmieden and Schott (2024); Gatti et al. (2024). GW abundance estimates suggest $\Omega_{\rm GW}\lesssim 10^{-12}$ at the peak.
V Conclusions
@@
In this paper, we have explored the astrophysical implications to a PBH population affected by the memory burden effect. Due to the associated halt in the evaporation process, light PBHs might live over cosmological timescales and constitute viable DM candidates, undergoing binary merging and forming new BHs of similar mass at much later cosmic times than their formation epochs. We argued that newly-formed BHs initially do not carry memory burden and can therefore emit energetic particles, in the form of cosmic rays, gamma rays, and neutrinos, which can be searched at dedicated facilities. We have estimated the integral spectrum in gamma rays from Galactic mergers and the differential spectrum of per-flavor neutrinos. Based on the current available measurements by IceCube, we have derived bounds on the PBH fraction. We found that, for a late memory burden onset, after the BH has lost roughly half of its initial mass, PBHs can account for the entirety of DM only in a small window $10^{9}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{10}$. Our findings are weakened if the memory burden sets in at parametrically earlier times. However, if the merger rate is increased by primordial non-Gaussianity, the mass range that is potentially excluded sensibly widens and our constraints strengthen. Future detectors will be able to probe even deeper into the remaining viable parameter space. Additional future directions include the careful study of the model with cosmic-ray and gamma-ray signatures from Galactic substructures.
Note Added. After the completion of this work, Ref. Dondarini et al. (2025) performed an in depth analysis of neutrinos, gamma-rays and early Universe signatures (CMB) resulting from the presented mechanism. We refer the interested reader to  Dondarini et al. (2025) for a detailed discussion of the existing constraints as well as the characterization of the duration of the semiclassical emission phase (given by $q$ in (1)) that allows $f_{\rm PBH}=1$ in the entire mass window. Our result in Fig. 3 is qualitatively consistent to theirs. Some minor discrepancies follow from the fact that Dondarini et al. (2025) adopted slightly different datasets and included the energy binning of the experimental points in the analysis.
@@ Acknowledgements.
We thank Gia Dvali for useful discussions and for reading the manuscript. We further thank Gwenael Giacinti for useful comments. Finally, we are grateful to Concha Gonzalez-Garcia for correcting a previous mistake in the detectability of the neutrino mass nature at IceCube and for further discussions. MZ is thankful to Alessandro Dondarini, Giulio Marino and Paolo Panci for useful ongoing discussion on related topics. We acknowledge support by the National Natural Science Foundation of China (NSFC) through the grant No. 12350610240 “Astrophysical Axion Laboratories”. LV acknowledges support by the Istituto Nazionale di Fisica Nucleare (INFN) through the “QGSKY” Iniziativa Specifica project. LV thanks the Tsung-Dao Lee Institute for its hospitality during the final stages of this work.